OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = 2/x^2+ 3 Is it g(x)=x^2 and F(x)=2/x^2+3?????? Please Help!!!

OpenStudy (anonymous):

\[f(x)= 2/x\] \[g(x) = x^2 +3\] So then you plug g(x) in for x in the function f(x)

OpenStudy (anonymous):

Thank you so much!